OpenStudy (anonymous):

how do you solve 7/2x - 5/3x= 22/3

OpenStudy (anonymous):

multiply everything by 6x: 21 - 10 = 44x x = 1/4

OpenStudy (anonymous):

factor out an x from the left side of the equation. x(7/2-5/3)=22/3 divide 22/3 with (7/2-5/3) you can multiply everything with 6 in order to get rid of the fractions (it will make the problem cleaner)